OpenStudy (anonymous):

A plant supervisor must apportion her 40-hour workweek between hours working on the assembly line and hours supervising the work of others. She is paid $12 per hour working and $15 per hour for supervising. If her earnings for a certain week are $504 how much time does she spend on each task?

OpenStudy (anonymous):

Lets call the number of hours working 'w' and the number of hours supervising 's'. From the question we know the total hours is 40: \[w+s=40\] We also know that total money earned is $540: \[12w+15s=540\] We can solve this by solving the first equation for w: \[w=40-s\] and substituting this into the second equation: \[12(40-s)+15s=540\]\[480-12s+15s=540\]\[3s+480=540\]\[s=20 hours\] We can now use our first equation to find that w=20. She supervises for 20 hours and works for 20 hours. We can check the answer with out equation: \[12w+15s=540\]\[12(20)+15(20)=540\]\[540 =540\]

OpenStudy (anonymous):

Thanks man!

OpenStudy (anonymous):

Aaaaannnd I just realized I did the whole question with $540 and you said $504.......the steps are right though. This is why I shouldn't do math at 1:30 in the morning.

OpenStudy (anonymous):

Fixed for 504" Same steps but you get 8 hours supervising ant 32 working. And the check works with the new numbers.